4mt 1. 3 m Find - Diameter a - Stress at bar 1&2 at pinc C 2 m 0 В. Pinat B roller at A Safety factor=3

Instead of diameter at pin C, calculate diameter at pin B
Answer #1

I'VE CALCULATED FOR STEEL AS A MATERIAL FOR PIN. BUT IF YOU HAVE OTHER METAL YOU CAN JUST REPLACE THE SIGMA VALUE IN LAST EQUATION TO GET ANSWER.

da tant(/2 = 63.43 * O = tem [34]t ton [2/4 B 2 Poby 63.430 3 RBT F-10kW Lanus egn at points F2 6314 10 Sin 63.4 F2 FI siho8.94 (0526:56 = KN Lanni eque at point A or tem-[34] = 3686 Be two.[4/30 = 53:13 2 Sins313 Rz = R3 ļ sin2rool sin 36.86 = -

PLEASE DO POST COMMENTS IF YOU HAVE ANY DOUBT. PRESS THUMBS UP IF YOU LIKE THE ANSWER AND THE EFFORT.

THANKS